OpenStudy (anonymous):

Solve and show your work? 4/7 x 1/8 =__?

OpenStudy (ghazi):

\[\frac{ 4 }{ 7 }*\frac{ 1 }{ 8 }=\frac{ 1 }{ 14 }\]

OpenStudy (ghazi):